# 3D-MVP: 3D Multiview Pretraining for Robotic Manipulation

### v1

@@ 4.4 Ablation Studies
Does the improvement come from the change of network architecture? In order to do the pretraining on RVT, we made some architecture changes as described in Sec. 3.2. To validate that the performance boost comes from the pretraining instead of the network architecture, we finetune our approach without pretraining and find the performance is 62.9, similar to the original RVT [15].
@@
Should we pretrain on object or room-level data? The choice of pretraining datasets are typically critical for self-supervised learning. In our experiments, we mainly use Objaverse [11], which is a object-centric 3D datasets. Since we are mainly evaluated on tabletop manipulation, we also try room-level 3D datasets such as 3D-FRONT [13]. We conduct the experiments on RLBench, and observe pretraining on 3D-FRONT only boosts the performance mildly from 62.9 to 63.6. In comparison, pretraining on Objaverse boosts the performance to 67.6. This suggests that the diversity and scale of the pretraining data are important for learning generalizable representations.
@@
Does more pretraining data help? We sampled a subset of Objaverse with 18K objects and called it Objaverse (small). The full Objaverse dataset we pretrain on has 200K objects. When we pretrain the encoder with Objaverse (small), we get a 65.3 mean success rate, which is worse than using Objaverse (full). This suggests that a larger dataset of pretraining helps performance in the downstream task.
@@
Figure 4: Pretraining MAE on only RLBench scenes leads poor generalization performance. (Left): MAE reconstruction results on unseen RLBench renderings. (Right): MAE reconstruction results on Objaverse renderings. While the reconstruction is reasonable on RLBench unseen renderings, it overfits to RLBench and does not learn a general representation.
@@
Can we pretrain on RLBench? The proposed 3D pretraining is self-supervised and only requires the 3D point cloud. Therefore, an interesting question is: can we just pretrain on the manipulation dataset (e.g. RLBench)? We extract the RLBench point cloud and build a pretraining dataset. After the pretraining, we finetune the model on training demonstrations as usual. We find it can achieve an average success rate of 67.5 on RLBench test set, which is comparable to Objaverse pretraining. However, the pretrained model suffers from generalization issues. As is shown in Figure 4, the encoder has overfitted to RLBench, and does not work on other environments such as COLOSSEUM.
@@
Masking strategy. For the masking strategy, we observe that masking RGB channels performs better than masking all channels. We hypothesize that the network finds it very challenging to reconstruct all the channels and is unable to learn better visual representations. We view our findings as similar to He et al. [19], who find that the benefits of pretraining diminish if the pretraining task becomes too difficult, like when the masking ratio becomes very high (> 80%).
@@ 5 Conclusion
Social impacts. The development of more generalized and robust robotic manipulation systems enabled by 3D-aware pretraining as proposed in our work has the potential for significant societal impact. On the positive side, such systems could automate many repetitive or dangerous manual labor tasks, improving worker safety and productivity. Assisting robots that can reliably manipulate objects in unstructured environments could also improve quality of life for the elderly and people with disabilities. However, the increased automation from these systems may also displace some jobs, disproportionately impacting workers with lower levels of education and technical skills. It will be important to enact policies to help retrain and transition affected workers to new roles. Additionally, the datasets used for pretraining these models, like Objaverse, may encode biases that could be reflected in the downstream robotic system’s behavior. Care must be taken to audit the data and model behavior to mitigate potential discriminatory impacts.
Acknowledgment. We thank Ishika Singh and Jiafei Duan for their help with the COLOSSEUM benchmark.

### v2

@@ 4.4 Ablation Studies
Does the performance boost come from pretraining or network architecture? We first validate whether the performance boost comes from architecture change in Sec. 3.2 or the pretraining itself. we finetune our model from scratch and find the performance is 62.9, similar to the original RVT [16]. It indicates the performance boost comes from our proposed 3D pretraining.
@@
Should we pretrain on object or room-level data? The choice of pretraining datasets are typically critical for self-supervised learning [9]. In our experiments, we mainly use Objaverse [10], which is a object-centric 3D datasets. Since we are mainly evaluated on tabletop manipulation, we also try room-level 3D datasets such as 3D-FRONT [14]. We conduct the experiments on RLBench, and observe pretraining on 3D-FRONT only boosts the performance mildly from 62.9 to 63.6. In comparison, pretraining on Objaverse boosts the performance to 67.6. We believe it is because 3D-FRONT has limited diversity and scale of objects and rooms compared with Objaverse. And the diversity and scale of the pretraining data are one of the most important factors for learning generalizable representations.
@@
Does more pretraining data help? To validate it, we sampled 18K objects from Objaverse and called it Objaverse (small). The full Objaverse dataset we use has 200K objects. When we pretrain the encoder with Objaverse (small), we get a 65.3 mean success rate, which is worse than using Objaverse (full). This suggests that a larger dataset of pretraining helps performance in the downstream task. If we can scale the training data to larger 3D model collections such as the full Objaverse-XL [11], that might improve the performance further.
@@
Figure 5: Pretraining MAE on RLBench scenes leads poor generalization performance. (Left): MAE reconstruction results on unseen RLBench renderings. (Right): MAE reconstruction results on Objaverse renderings. While the reconstruction is reasonable on RLBench unseen renderings, it overfits to RLBench and does not learn a general representation.
@@
Can we pretrain on RLBench? Instead of relying on a different large-scale pretraining dataset such as Objaverse, we investigate whether we can just pretrain on RLBench point clouds, since the proposed 3D pretraining is self-supervised and only requires the 3D point cloud. We extract the RLBench point cloud and build a pretraining dataset. After the pretraining, we finetune the model on training demonstrations as usual. We find it can achieve an average success rate of 67.5 on RLBench test set, which is comparable to Objaverse. However, the pretrained model suffers from generalization issues. As is shown in Figure 5, the encoder has overfitted to RLBench, and does not work on other environments such as COLOSSEUM.
@@
Masking strategy. We also observe that masking RGB channels performs better than masking all channels. We hypothesize that the network finds it very challenging to reconstruct all the channels and is unable to learn better visual representations. We view our findings as similar to He et al. [22], who find that the benefits of pretraining diminish if the pretraining task becomes too difficult, like when the masking ratio becomes very high ($>$80%).
@@ 5 Conclusion
Social impacts. The development of more generalized and robust robotic manipulation systems enabled by 3D-MVP has the potential for significant societal impact. On the positive side, such systems could automate many repetitive or dangerous manual labor tasks, improving worker safety and productivity. Assisting robots could also improve quality of life for the elderly and people with disabilities. However, the increased automation from these systems may also displace some jobs, disproportionately impacting workers with lower levels of education and technical skills. Additionally, the datasets used for pretraining these models, like Objaverse, may encode biases that could be reflected in the downstream robotic system’s behavior.
Acknowledgment. We thank Ishika Singh and Jiafei Duan for their help with COLOSSEUM, and Jesse Zhang for his help with some baseline results.